mix a solution that is 40% acid with a solution that is 100% water to make 16L of a solution that is 10% acid. How much of each solution should you use?
Answer by ptaylor(2198)

mix a solution that is 40% acid with a solution that is 100% water to make 16L of a solution that is 10% acid. How much of each solution should you use?
Let x=amount of 40% acid solution
Then 16-x=amount of 100% water.
Now we know the the amount of Pure acid in the 40% solution (.40x) plus the amount of pure acid in the 100% water solution (0) equals the amount of pure acid in the final solution (16)(.10). So, our equation to solve is:
.40x=(16)(.10)
.40x=1.6
x=4 liters of 40% solution
16-x=16-4=12 liters of pure water.
ck
(.4)(4)=16(.10)
1.6=1.6

Hope this helps----ptaylor